OpenStudy (aaronandyson):

The age of a father is twice the square of the age of his son.Eight years hence,the age of the father will be four years more than three times the age of the son.Find their present age.

OpenStudy (aaronandyson):

@sleepyjess

OpenStudy (aaronandyson):

@Directrix

OpenStudy (anonymous):

Let f and s be the current age of the father and the son respectively. Solve the following two equations for f and s: {f = 2 s^2, (f + 8) = 3 (s + 8) + 4} f = 32 and s = 4

OpenStudy (aaronandyson):

Can you explain it better?

OpenStudy (anonymous):

Is it the case that you do not see where the equations came from?

OpenStudy (aaronandyson):

no even I got this equation 2x^2 - 3x -20 = 0 I dont know what next now.

OpenStudy (anonymous):

The age of a father is twice the square of the age of his son. f = 2 s^2 32 = 2*4^2 =2*16=32 , so the solutions fit the first equation. Eight years hence is 8 years in the future from now. Dad and the Son will be f+8 and s+8 years old then. four years more than three times the age of the son relates to the numeric ages in the future.

OpenStudy (aaronandyson):

pfff nvm

OpenStudy (anonymous):

(f + 8) = 3 (s + 8) + 4 32+8 = 3 (4 + 8) + 4 40 = 36 +4 40 = 40 The age numbers satisfy the second equation.

OpenStudy (aaronandyson):

Thanks but no thanks.

OpenStudy (anonymous):

The difficulty with these age problems is converting the English language problem statements to algebra.

OpenStudy (aaronandyson):

Anyways Thanks

OpenStudy (anonymous):

Your welcome.